3x + 2y= -2
15x +6y= 12


answer:( 3, -11/2)

need a full break down of how my book came up with these solutions

Answer by checkley77(12569)

3x + 2y= -2 multipy this equation by -3 & add them.
15x +6y= 12
-9x-6y=6 [this equation is the 3x+2y=-2 equation multiplied by -3]
----------------------
6x=18 this equation is the sum of the 2 equations above.
x=18/6
x=3 ans.
3*3+2y=-2
9+2y=-2
2y=-2-9
2y=-11
y=-11/2
y=-5.5 ans.